# library.bib

### v1

@@ 5.1 Limitations
We apply random search with a comparatively low evaluation budget as a tuning strategy for our benchmark, which is not necessarily the best option for some of our slightly higher-dimensional search spaces. In particular for larger search spaces, Bayesian optimization [garnett2023bayesoptbook] could be more appropriate. However, we note that we greatly exceed the amount of tuning performed by many of the existing benchmarks in the literature (see Section 2). We distinguish between the ‘Cox’ and ‘AFT’ forms for XGBoost and treat them as separate models for the purposes of our benchmark, as the difference in objective represents very different model assumptions. Analogously, it could be argued that MBST should also be split in a similar fashion based on the family parameter. Furthermore, our use of the AFT model prioritizes prediction over interpretation and could be simplified by splitting the tuned model into one model for each functional form (“Weibull”, “log-normal”, “log-logistic”), which would yield more interpretable results. Finally, since our focus lies on generalizability to low-dimensional, right-censored settings, our results will not generalize to more complex settings. However, extension to left-censoring, competing risks or other more complex endpoints first necessitates more comprehensive support by models and their implementations as well as adaptation of available evaluation measures to these scenarios. The number of datasets included could be extended, but still exceeds that of the vast majority of previous benchmarks.
@@ 5.2 Conclusions and Future Work
Our results demonstrate that classical statistical methods, such as CPH and AFT, can significantly outperform complex ML algorithms in predictive survival tasks. While it is possible to fine-tune and achieve better predictive accuracy using ML methods in individual cases (see Appendix G), our results indicate that across a range of low-dimensional tasks, this is not the case in aggregate. We therefore recommend practitioners to start with these conceptually and computationally simpler methods, and evaluate whether the additional computational cost and loss of interpretability is appropriate for their needs. To improve upon our results, future work might employ alternative tuning strategies such as Bayesian Optimization or different model configurations with adjusted hyperparameter spaces. Additionally, an expansion of this benchmark with a wider range of settings would be beneficial, such that results can be generalized further pending the corresponding software support.
Acknowledgments and Disclosure of Funding
We are grateful for the package maintainers’ time supporting this effort regarding the model configurations used in this benchmark.
[…]
@@ Appendix A Literature Review
The experiments carried out in this paper fall into this category. Only two prior experiments could be found that neutrally benchmarked more than one ML model class on low-dimensional data. Kattan (2003) [kattan2003] benchmarked tree-based models, ANNs and CPH with Harrell’s C-index across three datasets with varying censoring proportions. The models are compared for significant differences by repeating the experiments up to 50 times with different seeds thus allowing for different hyper-parameter configurations and folds in cross-validation. Boxplots across all replications indicate that no machine learning model outperformed the CPH. Zhang et al. (2021) [zhang2021survbenchmark] compare classical and ML methods, taking into account feasibility and computational efficiency for various tasks in the biomedical field. Methods are evaluated on six clinical and 16 omics datasets using 11 metrics, including time-dependent AUC, Brier score and multiple variations of the C-index. However, methods were applied with specific hyperparameter sets without tuning, thereby limiting the generalizability of their results.
@@
Herrmann et al. (2020) [herrmann2020] performed a large-scale benchmark experiment of survival models on multi-omics high-dimensional data. Models fall into the following groups: Penalised regression, GBMs, and RSFs. Comparisons are made with Uno’s C and the Integrated Survival Brier Score (ISBS). The ISBS for all models overlapped with the Kaplan-Meier baseline though all C-indices were significantly higher than the baseline — however it is not stated how the standard errors for the confidence intervals were derived nor is it stated in the paper if multiple testing correction is applied. The authors also note that their results should be treated with caution due to the small performance differences and high variability. Spooner et al. (2020) [spooner2020] also compared machine learning models on high-dimensional data. In this study GBMs, RSFs, CPH and some extensions thereof were compared. Models were evaluated by Harrell’s C-index only. The results indicated that all models outperformed CPH when no additional feature selection was used but that there were no significant differences when feature selection was applied to the Cox model. There were few significant statistical differences between models. Wissel et al. (2023) [wissel2023systematic] provide a systematic comparison of multi-omics cancer survival models comparing eight DL methods, RSFs and CPH. They primarily focus on the noise-resistance of these models for high-dimensional settings, finding a general lack thereof when evaluating on Antolini’s C-index and ISBS.
@@
Moghimi-dehkordi et al. (2008) [moghimi-dehkordi2008] compare CPH to AFT models with various distributions. Out-of-sample measures for comparison are not provided though the AIC produced by the CPH is far higher (and therefore inferior) than those of the parametric models. Model inspection demonstrated that all models provided similar (non-significantly different) confidence intervals for hazard ratios. Georgousopoulou et al. (2015) [georgousopoulou2015] compared the CPH to a Weibull and Exponential AFT model. Again no out-of-sample measures were utilized, models were compared by the Cox-Snell residuals and the Bayesian Information Criterion (BIC). Similarly to Moghimi-dehkordi et al., hazard ratios produced from all three models were nearly identical. The authors claim the CPH is inferior to the parametric models though only graphical comparisons are included. Zare et al. (2015) [zare2015] provide another comparison of the CPH to AFT models, using Cox-Snell residuals and AIC as their measures of comparison. Similarly to the previous studies the Cox model has the highest AIC though the plotted Cox-Snell residuals are very similar. The authors acknowledge no significant differences between the model classes and instead conclude that AFT is a useful and more interpretable alternative. No significant differences were found between the different AFT parameterizations. Dirick et al. (2017) [dirick2017] make use of a financial setting to compare the CPH, AFT, flexible Cox models using splines to model the hazard, and mixture cure models. The authors compare the models using a time-dependent AUC, the mean squared error (MSE), and the mean absolute error (MAE). Survival times are generated from the CPH with a deterministic composition using quantiles chosen to minimize the MSE and MAE, and it is not clear if this is performed in an unbiased nested resampling manner or after predictions are made. By averaging the ranking of model performance the authors conclude that CPH with penalized splines outperformed the other models with respect to the chosen metrics.
Habibi et al. (2018) [habibi2018] performed another experiment on PH and AFT models. Models were again compared exclusively by the AIC with the PH having the highest result and log-normal AFT the lowest; differences between AFT models were non-significant. Confidence intervals for hazard ratios were similar (non-significantly different) for all models.
@@
Luxhoj and Shyur (1997) [luxhoj1997] benchmarked neural networks against CPH in the engineering field of reliability analysis. The baseline hazard of the Cox model is modeled by splines with a single knot. The models are compared using the mean squared error on a validation set of sample size 40, of these 40 there are only 9 unique failure times that are used for model testing. Insufficient information is provided to determine the architecture or training procedure of the neural networks compared. The MSE difference between the Cox and ANN was 0.003, which is highly unlikely to be a significant difference on a test set of only 40 observations with nine observed events. Ohno-Machado (1997) [ohno-machado1997] also compared CPH to ANNs. Several Cox models were fit with automated variable selection by backwards elimination. For each model, survival curves were predicted and for a given patient they were considered dead at a particular time point if the predicted survival curve at the time is less than the “arbitrary” [ohno-machado1997] probability of $0.5$. The Cox models were compared to a single hidden layer ANN. This model made probabilistic predictions of death in four time-intervals that were within the predicted time of the Cox models. The probabilistic predictions from both models were compared with the AUC and its corresponding ROC. No significant differences in performance were found between the two models. Goli et al. (2016) [goli2016a] provide a comprehensive comparison of support vector machine models with CPH as a reference class (Kaplan-Meier is not included). Models are compared against the C-index and log-rank test, though it is unstated which C-index is utilized. No model outperformed CPH with respect to the chosen C-index. Jaeger et al. (2019) [jaeger2024] compare a multiple variations of a novel implementation of oblique RSFs (“aorsf”) to the previous implementation, as well as other RSFs, GBMs, penalized CPH, and ANNs. They compare methods on 21 datasets, including low- and high-dimensional settings. Results are analysed using post-hoc Bayesian ROPE and evaluated using the Index of Prediction Accuracy (IPA) [kattan2018indexprediction] based on the ISBS and time-dependent C-index [blanche2013estimatingcomparing]. They present results relative to the method performing best in their benchmark (“aorsf-fast” for both measures), with GBMs among the lowest-performing methods. Only minimal tuning was conducted.
@@ A.1 Surveys of Survival Models
The final class of papers do not perform empirical benchmark experiments but instead survey/review available survival models. These are therefore only discussed very briefly. Ohno-Machado (2001) [ohno-machado2001] provide an overview to models available for survival analysis from non-parametric estimators and classical models to neural networks. The review highlights useful applications of the models and their respective limitations. In particular their Table 1 clearly states advantages and disadvantages of Cox models versus ANNs. Patel et al. (2006) [patel2006] compare proportional hazards and accelerated failure time models. This comparison is primarily theoretical and based on model properties, no analytical comparison with measures is provided though comparisons of predicted median survival times are compared to those from a Kaplan-Meier estimator. The authors conclude that AFT models should be considered more often due to simpler interpretation. Wang et al. (2019) [wang2019] provide a review of survival analysis models and measures that is strongly recommended here as a precise and comprehensive introduction to the field of machine learning in survival analysis. The authors provide strong arguments for comparing classical models against one another, though this is not extended to the machine learning setting. More mathematical detail is provided to the classical setting however a clear and detailed overview is still provided for machine learning models. Some attention is also given to the more complex cases of competing risks and multiple events. Lee and Lim (2019) [lee2019] provide a short but concise overview to survival analysis models with an emphasis on genetic data and implementation in R. Their review covers classical models, penalisation, and many machine learning models. They provide a clear, practical illustration (but no full benchmark experiment) comparing the models on a real dataset against Harrell’s C. No model outperforms CPH. Wiegrebe et al. (2024) [wiegrebe2024deeplearning] provide a comprehensive overview of deep learning methods for survival analysis and compare methods based on their capabilities regarding various common challenges in survival analysis such as time-varying features, competing events, different censoring types, dimensionality, modality, and interpretability. Their extensive comparison is also available as a web-based interactive table.
@@ Appendix B Datasets
Table 2 lists datasets included in the benchmark along with their sources and common descriptive statistics. Table 3 lists the licenses declared by the source packages from which the dataset is taken.
@@
aids.id [dataaidsid] 60 1 4 467 5 188 JM [pkgjm]
aids2 [pkgnnet] 38 1 3 2814 4 1733 MASS [pkgnnet]
bladder0 [databladder0] 48 0 3 397 3 206 frailtyHL [pkgfrailtyhl]
CarpenterFdaData [carpenter2002] 36 15 11 408 26 262 simPH
channing [klein2003] 62 1 1 458 2 176 KMsurv
child [pkgeha] 79 1 3 26574 4 5616 eha [pkgeha]
colrec [pkgrelsurv] 17 3 2 5578 13 4602 relsurv [pkgrelsurv]
cost [datacost] 22 3 10 518 13 404 pec [pkgpec]
dataFTR [dataftrstr] 86 0 2 2206 2 300 RISCA [pkgrisca]
dataSTR [dataftrstr] 82 0 4 546 4 101 RISCA
e1684 [datae1684] 31 1 2 284 3 196 smcure [pkgsmcure]
flchain [dataflchain] 72 4 3 7871 7 1082 survival
gbsg [katzman2018] 43 3 4 2232 7 1267 pycox [pkgpycox]
grace [dataapplied] 68 4 2 1000 6 324 mlr3proba [pkgmlr3proba]
hdfail [pkgfrailtysurv] 94 1 4 52422 5 2885 frailtySurv [pkgfrailtysurv]
kidtran [klein2003] 84 1 3 863 4 140 KMsurv
liver [dataliver] 40 1 1 488 2 292 joineR [pkgjoiner]
lung [datalung] 28 5 3 167 8 120 survival
metabric [katzman2018] 42 5 4 1903 9 1103 pycox
mgus [datamgus] 6 6 1 176 7 165 survival
nafld1 [datanafld1] 92 4 1 12446 5 1018 survival
nwtco [datanwtco] 86 1 2 4028 3 571 survival
ova [dataova] 26 1 4 358 5 266 dynpred
patient [datapatient] 79 2 5 1985 7 416 pammtools [pkgpammtools]
rdata [pkgrelsurv] 47 1 3 1040 4 547 relsurv [pkgrelsurv]
std [klein2003] 60 3 18 877 21 347 KMsurv
support [katzman2018] 32 10 4 8873 14 2705 pycox
tumor [pkgpammtools] 52 1 6 776 7 375 pammtools
uis [datauis] 19 7 5 575 12 464 quantreg [pkgquantreg]
veteran [kalbfleisch2011] 7 3 3 137 6 128 survival
wbc1 [datawbc] 43 2 0 190 4 109 dynpred
whas [dataapplied] 48 3 6 481 9 249 mlr3proba
@@
GPL (>= 2) jm, eha, pec, RISCA, dynpred, quantreg
GPL (>= 3) KMsurv
@@
Appendix C Models
LABEL:tab:models lists all the compared models, along with their software, version numbers, and prediction types. The horizontal lines separate the models into different groups: Baseline, Classical, Random Survival Forests (RSFs), Gradient Boosting Machines (GBMs) and Support Vector Machines (SVMs).

### v2

@@ 5.1 Limitations
Our use of the AFT model prioritizes prediction over interpretation and could be simplified by splitting the tuned model into one model for each functional form (“Weibull”, “log-normal”, “log-logistic”), which would yield more interpretable results. Since our focus lies on generalizability to low-dimensional, right-censored settings, our results will not generalize to more complex settings. However, extension to left-censoring, competing risks, or other more complex endpoints first necessitates more comprehensive support by models and their implementations, as well as adaptation of available evaluation measures to these scenarios. The number of datasets included could be extended, but still exceeds that of the vast majority of previous benchmarks.
@@ 5.2 Conclusions and Future Work
Our results demonstrate that classical statistical methods, such as CPH and AFT, can in some settings significantly outperform complex ML algorithms in predictive survival tasks. While it is possible to fine-tune and achieve better predictive accuracy using ML methods in individual cases (see Appendix F), our results indicate that across a range of low-dimensional tasks, this is not the case in aggregate. We therefore recommend practitioners to start with these conceptually and computationally simpler methods, and evaluate whether the additional computational cost and loss of interpretability are appropriate for their needs. To improve upon our results, expanding this benchmark with a wider range of settings would be beneficial, allowing results to be generalized further, pending the corresponding software support.
@@ Appendix A Literature Review
The experiments carried out in this paper fall into this category. Only two prior experiments could be found that neutrally benchmarked more than one ML model class on low-dimensional data. Kattan (2003) benchmarked tree-based models, ANNs and CPH with Harrell’s C-index across three datasets with varying censoring proportions. The models are compared for significant differences by repeating the experiments up to 50 times with different seeds thus allowing for different hyper-parameter configurations and folds in cross-validation. Boxplots across all replications indicate that no machine learning model outperformed the CPH. Zhang et al. (2021) compare classical and ML methods, taking into account feasibility and computational efficiency for various tasks in the biomedical field. Methods are evaluated on six clinical and 16 omics datasets using 11 metrics, including time-dependent AUC, Brier score and multiple variations of the C-index. However, methods were applied with specific hyperparameter sets without tuning, thereby limiting the generalizability of their results.
@@
Herrmann et al. (2021) performed a large-scale benchmark experiment of survival models on multi-omics high-dimensional data. Models fall into the following groups: Penalised regression, GBMs, and RSFs. Comparisons are made with Uno’s C and the Integrated Survival Brier Score (ISBS). The ISBS for all models overlapped with the Kaplan-Meier baseline though all C-indices were significantly higher than the baseline — however it is not stated how the standard errors for the confidence intervals were derived nor is it stated in the paper if multiple testing correction is applied. The authors also note that their results should be treated with caution due to the small performance differences and high variability. Spooner et al. (2020) also compared machine learning models on high-dimensional data. In this study GBMs, RSFs, CPH and some extensions thereof were compared. Models were evaluated by Harrell’s C-index only. The results indicated that all models outperformed CPH when no additional feature selection was used but that there were no significant differences when feature selection was applied to the Cox model. There were few significant statistical differences between models. Wissel et al. (2023) provide a systematic comparison of multi-omics cancer survival models comparing eight DL methods, RSFs and CPH. They primarily focus on the noise-resistance of these models for high-dimensional settings, finding a general lack thereof when evaluating on Antolini’s C-index and ISBS.
@@
Moghimi-Dehkordi et al. (2008) compare CPH to AFT models with various distributions. Out-of-sample measures for comparison are not provided though the AIC produced by the CPH is far higher (and therefore inferior) than those of the parametric models. Model inspection demonstrated that all models provided similar (non-significantly different) confidence intervals for hazard ratios. Georgousopoulou et al. (2015) compared the CPH to a Weibull and Exponential AFT model. Again no out-of-sample measures were utilized, models were compared by the Cox-Snell residuals and the Bayesian Information Criterion (BIC). Similarly to Moghimi-dehkordi et al., hazard ratios produced from all three models were nearly identical. The authors claim the CPH is inferior to the parametric models though only graphical comparisons are included. Zare et al. (2015) provide another comparison of the CPH to AFT models, using Cox-Snell residuals and AIC as their measures of comparison. Similarly to the previous studies the Cox model has the highest AIC though the plotted Cox-Snell residuals are very similar. The authors acknowledge no significant differences between the model classes and instead conclude that AFT is a useful and more interpretable alternative. No significant differences were found between the different AFT parameterizations. Dirick et al. (2017) make use of a financial setting to compare the CPH, AFT, flexible Cox models using splines to model the hazard, and mixture cure models. The authors compare the models using a time-dependent AUC, the mean squared error (MSE), and the mean absolute error (MAE). Survival times are generated from the CPH with a deterministic composition using quantiles chosen to minimize the MSE and MAE, and it is not clear if this is performed in an unbiased nested resampling manner or after predictions are made. By averaging the ranking of model performance the authors conclude that CPH with penalized splines outperformed the other models with respect to the chosen metrics.
Habibi et al. (2018) performed another experiment on PH and AFT models. Models were again compared exclusively by the AIC with the PH having the highest result and log-normal AFT the lowest; differences between AFT models were non-significant. Confidence intervals for hazard ratios were similar (non-significantly different) for all models.
@@
Luxhoj and Shyur (1997) benchmarked neural networks against CPH in the engineering field of reliability analysis. The baseline hazard of the Cox model is modeled by splines with a single knot. The models are compared using the mean squared error on a validation set of sample size 40, of these 40 there are only 9 unique failure times that are used for model testing. Insufficient information is provided to determine the architecture or training procedure of the neural networks compared. The MSE difference between the Cox and ANN was 0.003, which is highly unlikely to be a significant difference on a test set of only 40 observations with nine observed events. Ohno-Machado (1997) also compared CPH to ANNs. Several Cox models were fit with automated variable selection by backwards elimination. For each model, survival curves were predicted and for a given patient they were considered dead at a particular time point if the predicted survival curve at the time is less than the “arbitrary”  Ohno-Machado (1997) probability of $0.5$. The Cox models were compared to a single hidden layer ANN. This model made probabilistic predictions of death in four time-intervals that were within the predicted time of the Cox models. The probabilistic predictions from both models were compared with the AUC and its corresponding ROC. No significant differences in performance were found between the two models. Goli et al. (2016) provide a comprehensive comparison of support vector machine models with CPH as a reference class (Kaplan-Meier is not included). Models are compared against the C-index and log-rank test, though it is unstated which C-index is utilized. No model outperformed CPH with respect to the chosen C-index. Jaeger et al. (2024) compare a multiple variations of a novel implementation of oblique RSFs (“aorsf”) to the previous implementation, as well as other RSFs, GBMs, penalized CPH, and ANNs. They compare methods on 21 datasets, including low- and high-dimensional settings. Results are analyzed using post-hoc Bayesian ROPE and evaluated using the Index of Prediction Accuracy (IPA) (Kattan and Gerds, 2018) based on the ISBS and time-dependent C-index (Blanche et al., 2013). They present results relative to the method performing best in their benchmark (“aorsf-fast” for both measures), with GBMs among the lowest-performing methods. Only minimal tuning was conducted.
@@ A.1 Surveys of Survival Models
The final class of papers do not perform empirical benchmark experiments but instead survey/review available survival models. These are therefore only discussed very briefly. Ohno-Machado (2001) provide an overview to models available for survival analysis from non-parametric estimators and classical models to neural networks. The review highlights useful applications of the models and their respective limitations. In particular their Table 1 clearly states advantages and disadvantages of Cox models versus ANNs. Patel et al. (2006) compare proportional hazards and accelerated failure time models. This comparison is primarily theoretical and based on model properties, no analytical comparison with measures is provided though comparisons of predicted median survival times are compared to those from a Kaplan-Meier estimator. The authors conclude that AFT models should be considered more often due to simpler interpretation. Wang et al. (2019) provide a review of survival analysis models and measures that is strongly recommended here as a precise and comprehensive introduction to the field of machine learning in survival analysis. The authors provide strong arguments for comparing classical models against one another, though this is not extended to the machine learning setting. More mathematical detail is provided to the classical setting however a clear and detailed overview is still provided for machine learning models. Some attention is also given to the more complex cases of competing risks and multiple events. Lee and Lim (2019) provide a short but concise overview to survival analysis models with an emphasis on genetic data and implementation in R. Their review covers classical models, penalisation, and many machine learning models. They provide a clear, practical illustration (but no full benchmark experiment) comparing the models on a real dataset against Harrell’s C. No model outperforms CPH. Wiegrebe et al. (2024) provide a comprehensive overview of deep learning methods for survival analysis and compare methods based on their capabilities regarding various common challenges in survival analysis such as time-varying features, competing events, different censoring types, dimensionality, modality, and interpretability. Their extensive comparison is also available as a web-based interactive table.
@@ Appendix B Datasets
Table˜2 lists datasets included in the benchmark along with their sources and common descriptive statistics. Table˜3 lists the licenses declared by the source packages from which the dataset is taken.
@@
aids.id (Carlin and Louis, 2018) 60 1 4 467 5 188 JM (Rizopoulos, 2010)
aids2 (Venables and Ripley, 2002) 38 1 3 2814 4 1733 MASS (Venables and Ripley, 2002)
bladder0 (Sylvester et al., 2006) 48 0 3 397 3 206 frailtyHL (Ha et al., 2019)
CarpenterFdaData (Carpenter, 2002) 36 15 11 408 26 262 simPH
cat_adoption (Kuhn, 2024) 37 14 4 2257 18 1434 modeldata (Kuhn, 2024)
channing (Klein and Moeschberger, 2003) 62 1 1 458 2 176 KMsurv
check_times (Kuhn, 2024) 1 22 0 13626 22 13523 modeldata
child (Broström, 2021) 79 1 3 26574 4 5616 eha (Broström, 2021)
colrec (M. Pohar and J. Stare, 2006) 17 3 2 5578 13 4602 relsurv (M. Pohar and J. Stare, 2006)
cost (Jørgensen et al., 1996) 22 3 10 518 13 404 pec (Mogensen et al., 2014)
dataFTR (Trébern-Launay et al., 2013) 86 0 2 2206 2 300 RISCA (Foucher et al., 2023)
dataSTR (Trébern-Launay et al., 2013) 82 0 4 546 4 101 RISCA
e1684 (Kirkwood et al., 1996) 31 1 2 284 3 196 smcure (Cai et al., 2012)
flchain (Dispenzieri et al., 2012) 72 4 3 7871 7 1082 survival
gbsg (Katzman et al., 2018) 43 3 4 2232 7 1267 pycox (Kvamme, 2018)
grace (Hosmer Jr et al., 2011) 68 4 2 1000 6 324 mlr3proba (Sonabend et al., 2021)
hdfail (Monaco et al., 2018) 94 1 4 52422 5 2885 frailtySurv (Monaco et al., 2018)
kidtran (Klein and Moeschberger, 2003) 84 1 3 863 4 140 KMsurv
liver (Andersen et al., 1993) 40 1 1 488 2 292 joineR (Williamson et al., 2008)
lung (Loprinzi et al., 1994) 28 5 3 167 8 120 survival
metabric (Katzman et al., 2018) 42 5 4 1903 9 1103 pycox
mgus (Kyle, 1993) 6 6 1 176 7 165 survival
nafld1 (Allen et al., 2018) 92 4 1 12446 5 1018 survival
nwtco (Breslow and Chatterjee, 1999) 86 1 2 4028 3 571 survival
ova (Van Houwelingen et al., 1989) 26 1 4 358 5 266 dynpred
rdata (M. Pohar and J. Stare, 2006) 47 1 3 1040 4 547 relsurv (M. Pohar and J. Stare, 2006)
std (Klein and Moeschberger, 2003) 60 3 18 877 21 347 KMsurv
support (Katzman et al., 2018) 32 10 4 8873 14 2705 pycox
tumor (Bender and Scheipl, 2018) 52 1 6 776 7 375 pammtools
uis (Hosmer et al., 2008) 19 7 5 575 12 464 quantreg (Koenker, 2021)
veteran (Kalbfleisch and Prentice, 2011) 7 3 3 137 6 128 survival
wa_churn (Kuhn, 2024) 73 8 10 7032 18 1869 modeldata
wbc1 (The Benelux C M L Study Group, 1998) 43 2 0 190 4 109 dynpred
whas (Hosmer Jr et al., 2011) 48 3 6 481 9 249 mlr3proba
@@
GPL ($\geq$ 2) jm, eha, pec, RISCA, dynpred, quantreg
GPL ($\geq$ 3) KMsurv
@@
Appendix C Models and Configurations
LABEL:tab:models lists all the compared algorithms, along with their R package name, and prediction types. The horizontal lines separate the models into different groups: (1) Baseline learners (2) Classical parametric models including penalized versions (3) Tree-based methods, including random survival forests and Trees (4) Boosting algorithms including gradient- and likelihood-boosting (5) Support Vector Machines (SVMs). LABEL:tab:config shows hyperparameter search spaces and non-default parameter values as well as common pre-processing requirements.